# Summaries as Centroids for Interpretable and Scalable Text Clustering

### v4

@@ Abstract
We introduce k-NLPmeans and k-LLMmeans, text-clustering variants of k-means that periodically replace numeric centroids with textual summaries. The key idea—summary-as-centroid—retains k-means assignments in embedding space while producing human-readable, auditable cluster prototypes. The method is LLM-optional: k-NLPmeans uses lightweight, deterministic summarizers, enabling offline, low-cost, and stable operation; k-LLMmeans is a drop-in upgrade that uses an LLM for summaries under a fixed per-iteration budget whose cost does not grow with dataset size. We also present a mini-batch extension for real-time clustering of streaming text. Across diverse datasets, embedding models, and summarization strategies, our approach consistently outperforms classical baselines and approaches the accuracy of recent LLM-based clustering—without extensive LLM calls. Finally, we provide a case study on sequential text streams and release a StackExchange-derived benchmark for evaluating streaming text clustering.
@@ 1 Introduction
Text clustering is a core problem in natural language processing (NLP), with applications in document organization, topic exploration, and information retrieval [55, 61]. A standard pipeline embeds documents into vectors [17, 54, 38, 46, 10, 29] and then groups them with a clustering algorithm [47]. Among these algorithms, k-means [36] remains ubiquitous, iteratively updating each centroid as the mean of its assigned points. While effective, this purely numerical averaging can blur contextual nuance present in the original texts [52]. Prior work has explored alternative centroid definitions and related objectives [27, 9, 31], yet these approaches remain anchored in vector space, which limits interpretability and can induce semantic drift between centroids and their underlying documents. This motivates centroids that are explicitly textual, aligning prototypes with human-interpretable summaries.
@@
We consider two families of summarizers: (1) Classical NLP such as centroid-based summarization [48], graph-based methods such as TextRank [37], and LSA-style techniques [14], which are fast and deterministic (yielding k-NLPmeans); and (2) LLM-based summaries that can capture richer context and filter secondary or noisy content [71, 50, 28] (yielding k-LLMmeans). In both variants, summaries are recomputed only at scheduled iterations, so the algorithms share the same summary-as-centroid mechanism and differ solely in the summarizer. While methods from (1) and (2) are standard in summarization, our periodic use of summary-derived centroids within the k-means loop is, to our knowledge, novel and central to the method’s interpretability.
@@
Each summarization step produces a concise, human-readable centroid that exposes evolving cluster semantics, simplifying debugging, validation, and labeling. The summarizer is modular: it can be LLM-free (k-NLPmeans) or LLM-assisted under a fixed per-iteration budget that does not grow with dataset size (k-LLMmeans). For large-scale and streaming settings, we adapt mini-batch k-means [56] by inserting summarization steps into the mini-batch update rule, enabling efficient online clustering while retaining interpretability.
@@
Recent work shows that LLMs can deliver strong unsupervised clustering [73, 20, 13, 64, 58, 62, 39]. However, many pipelines face two practical issues: (i) scalability, as they may rely on semi-supervision, iterative labeling, or a number of LLM calls that grows with dataset size; and (ii) opaque optimization, since they often combine prompts, greedy merges, and similarity thresholds without an explicit objective, making convergence behavior hard to analyze. Our k-LLMmeans addresses both: by summarizing per cluster at spaced iterations, we cap LLM usage independently of dataset size, and by keeping assignments and numeric updates in embedding space, we preserve the standard k-means objective between summary steps. In both k-NLPmeans and k-LLMmeans, if the summaries are weak, the procedure gracefully degrades to vanilla k-means with a suboptimal initialization, retaining its standard local-convergence behavior. Note that our method is LLM-optional: k-NLPmeans replaces the summarizer with classical NLP techniques, incurring no LLM usage.
@@ 2 Preliminaries: k-means for text clustering
$\min_{C_{1},C_{2},\dots,C_{k}}\sum_{j=1}^{k}\sum_{i\in[C_{j}]}\|\mathbf{x}_{i}-\boldsymbol{\mu}_{j}\|^{2},$ (1)
where $C_{j}$ denotes the set of embeddings assigned to cluster $j$, $C_{j}]=\{i|\mathbf{x}_{i}\in C_{j}\}$ denotes the set of embedding indices assigned to cluster $j$ and $\boldsymbol{\mu}_{j}$ is the cluster centroid, computed as the mean of the assigned embeddings:
$\boldsymbol{\mu}_{j}=\frac{1}{|C_{j}|}\sum_{i\in[C_{j}]}\mathbf{x}_{i}.$ (2)
Lloyd’s algorithm [35], the standard heuristic for k-means, alternates between assigning each document embedding $\mathbf{x}_{i}$ to its nearest centroid and recomputing each centroid as the mean of its assigned points. Repeating these two steps for $T$ iterations monotonically reduces the within‑cluster sum of squared distances, steering the procedure toward a (locally) optimal set of centroids. However, due to its sensitivity to initialization and the non-convex nature of its objective function, k-means does not guarantee convergence to the global optimum and can instead become trapped in local optima [36, 35]. Various strategies, such as k-means++ initialization and multiple restarts, have been proposed to mitigate these issues and improve the likelihood of achieving better clustering results [5].
@@
We enhance k-means for text clustering by periodically replacing the numerical centroid update with summarization steps that yield textual-summary–based centroids. The procedure (Algorithm 1 in Appendix D) is identical to k-means algorithm except that every $l$ iterations the mean update in (2) is replaced by the embedding of a cluster textual summary. At all other iterations, the standard update is used (see Figure 1 for an illustration with one summarization step). The summarizer can be instantiated with classical, deterministic methods (k-NLPmeans) or with LLM-based summaries (k-LLMmeans), and is not restricted to the specific techniques we discuss, any textual summarization operator can be dropped in.
@@
Our first variant uses classical extractive summarization to compute a textual prototype in place of the standard centroid update. Formally, we replace (2) with
$\boldsymbol{\mu}_{j}=\text{Embedding}\!\left(f^{(q)}_{\text{NLP}}(S_{j})\right),$ (3)
@@
  • Centroid-based summarization [48]: compute the centroid of the sentence embeddings for $S_{j}$; rank sentences by cosine similarity to this centroid; concatenate the top $q$ sentences (by rank) to form $f^{(q)}_{\text{NLP}}(S_{j})$.
  • Graph-based (TextRank) [37]: build a graph whose nodes are sentences in $S_{j}$ with edge weights given by pairwise cosine similarity of their embeddings; run a PageRank-style algorithm to score sentences; select and concatenate the top $q$.
  • LSA-style SVD in embedding space [14]: stack the sentence embeddings for $S_{j}$, apply singular value decomposition, score sentences by contribution to the leading components, and concatenate the top $q$.
The resulting summary text $f^{(q)}_{\text{NLP}}(S_{j})$ is then embedded in the same vector space as the documents to yield the new centroid $\boldsymbol{\mu}_{j}$.
@@
If instead we generate the summary with an LLM, we have k-LLMmeans. Formally, we replace (2) by
$\boldsymbol{\mu}_{j}=\text{Embedding}(f_{\text{LLM}}(p_{j}))\\$ (4)
where $p_{j}=\text{Prompt}\left(I,\{d_{z_{i}}|z_{i}\sim[C_{j}]\}_{i=1}^{m_{j}}\right)$, and $m_{j}=\min(m,|C_{j}|)$. Here, $z_{i}\sim[C_{j}]$ denotes a sampled index of the embeddings assigned to cluster $C_{j}$ (without repetitions) and $m$ is a parameter that represents the maximum number of sampled indices used to compute the cluster centroid $\boldsymbol{\mu}_{j}$. In simple terms, we update a cluster’s centroid by using the embedding of the response generated by an LLM when queried with a prompt containing a summarization instruction $I$ and a representative sample of documents from the cluster. Rather than providing all documents within the cluster as input, the LLM processes a representative sample as a context prompt. While incorporating the entire cluster is theoretically possible, it poses practical challenges due to prompt length limitations. Therefore, we propose selecting the sample cluster documents using a k-means++ sampling of the cluster embeddings. Our experiments demonstrate that this sampling process facilitates a more effective synthesis of the cluster’s content, leading to improved summaries and, consequently, more refined centroid updates. The instruction $I$ varies depending on the clustering task, but standard summarization prompts are generally sufficient. Figure 1 illustrates how k-LLMmeans with a single summarization step enhances the standard k-means algorithm.
@@ 3.3 Advantages of our approaches
Over advanced LLM-based clustering methods. Our approach offers three key advantages over more complex LLM-based clustering methods: (1) Optimization landscape. Grounded in the standard k-means objective, we inherit algorithmic convergence without relying on the fragile heuristics common in newer LLM‑driven methods. A poor summary simply leads the procedure back toward a regular k-means local optimum, whereas competing approaches hinge on stable, format‑specific LLM outputs. (2) Scalability. Unlike most state-of-the-art methods, whose LLM usage complexity grows with sample size [20, 13] or require fine-tuning [73]. k-NLPmeans requires no LLM calls and k-LLMmeans only performs $k$ LLM calls per summarization step, with even few summarization steps yielding substantial performance gains. (See Section 5.1). (3) Interpretability. Replacing numeric centroids with textual summaries turns each prototype into a concise, human-readable synopsis; practitioners can track how cluster semantics evolve over time without post-hoc labeling. This transparency extends naturally to our mini-batch variant, enabling real-time monitoring in streaming scenarios (see Figure 2 and Section 5.1).
@@ 4 Mini-batch k-NLPmeans and k-LLMmeans
Mini-batch k-means [56] is an efficient strategy for large-scale text clustering that processes small, randomly sampled mini-batches instead of the full dataset. This approach substantially reduces memory usage and computational cost, making it well suited for continuously generated text streams, such as those from social media, news, or customer feedback; where data must be clustered incrementally without full dataset access. Mini-batch k-means exhibits convergence properties comparable to standard k-means while offering superior scalability.
Although numerous streaming clustering methods that do not rely on LLMs have been studied [59, 3, 53, 2, 1, 45], only a few incorporate LLMs [62, 39]. Moreover, existing offline LLM-based clustering approaches face scalability issues, highlighting the need for scalable summary-based clustering in an online setting. To address this, we introduce mini-batch k-NLPmeans and k-LLMmeans, which directly extend mini-batch k-means by inserting summarization steps into the mini-batch update.
@@ 5 Static experiments
We use four benchmark datasets that span diverse domains and classification granularities: Bank77 [12], CLINC [33], GoEmo [16] and MASSIVE (domain and intent) [21]. See Appendix A.1 for detailed descriptions of the datasets. We evaluate our algorithms on each of the four datasets using the known number of clusters and performing 120 centroid-update iterations. We calculate two variants of our algorithms differing in the number of summarization steps. The single variant uses a single summarization step ($l=60$), while the multiple variant performs five summarization steps ($l=20$). To demonstrate the robustness of our approach, we compute embeddings with models: DistilBERT[54], e5-large[65], S-BERT[52], and text-embedding-3-small[44]. For k-NLPmeans we evaluate using TextRank, Centroid and LSA summarization methods mentioned in Section 3.1, with $q=5$. For the LLM component of k-LLMmeans, we use GPT-3.5-turbo[43], GPT-4o[26], Llama-3.3[22], Claude-3.7[4] and DeepSeek-V3[34]. For the instruction task $I$, we employ a simple summarization prompt depending on the task. For example, for Bank77 we use the prompt: “The following is a cluster of online banking questions. Write a single question that represents the cluster concisely.” We explore the effect of prompt size by summarizing using all cluster documents as input, and a few-shot (FS) variant where each prompt includes only $m=10$ randomly selected documents instead of the full cluster. For each setting we run five different seeds.
As traditional clustering baselines, we consider k-means, k-medoids [31], spectral clustering [41], agglomerative clustering [30], and Gaussian Mixture Models (GMM) [15]. We use ground-truth number of clusters and ten different seeds. We also include BERTopic [23] as a strong topic-modeling baseline. In addition, we report results obtained by applying our proposed methods to the document embeddings produced by BERTopic. For advanced LLM-based methods, we compare with results obtained by Feng et al. [20] on ClusterLLM [73], IDAS [13], and LLMEdgeRefine [20]. All k-means based methods are initialized with k-means++.

### v5

@@ Abstract
We introduce k-NLPmeans and k-LLMmeans, text-clustering variants of k-means that periodically replace numeric centroids with textual summaries. The key idea, summary-as-centroid, retains k-means assignments in embedding space while producing human-readable, auditable cluster prototypes. The method is LLM-optional: k-NLPmeans uses lightweight, deterministic summarizers, enabling offline, low-cost, and stable operation; k-LLMmeans is a drop-in upgrade that uses an LLM for summaries under a fixed per-iteration budget whose cost does not grow with dataset size. We also present a mini-batch extension for real-time clustering of streaming text. Across diverse datasets, embedding models, and summarization strategies, our approach consistently outperforms classical baselines and approaches the accuracy of recent LLM-based clustering without extensive LLM calls. Finally, we provide a case study on sequential text streams and release a StackExchange-derived benchmark for evaluating streaming text clustering.
@@ 1 Introduction
Text clustering is a core problem in natural language processing (NLP), with applications in document organization, topic exploration, and information retrieval (Schütze et al., 2008; Steinbach, 2000). A standard pipeline embeds documents into vectors (Devlin, 2018; Sanh, 2019; Mikolov, 2013; Pennington et al., 2014; Brown et al., 2020; Jin et al., 2023) and then groups them with a clustering algorithm (Petukhova et al., 2025). Among these algorithms, k-means (MacQueen, 1967) remains ubiquitous, iteratively updating each centroid as the mean of its assigned points. While effective, this purely numerical averaging can blur contextual nuance present in the original texts (Reimers and Gurevych, 2019). Prior work has explored alternative centroid definitions and related objectives (Jain and Dubes, 1988; Bradley et al., 1996; Kaufman and Rousseeuw, 2008), yet these approaches remain anchored in vector space, which limits interpretability and can induce semantic drift between centroids and their underlying documents. This motivates centroids that are explicitly textual, aligning prototypes with human-interpretable summaries.
@@
We consider two families of summarizers: (1) Classical NLP such as centroid-based summarization, graph-based methods such as TextRank , and LSA-style techniques (Radev et al., 2004; Mihalcea and Tarau, 2004; Deerwester et al., 1990), which are fast and deterministic (yielding k-NLPmeans); and (2) LLM-based summaries that can capture richer context and filter secondary or noisy content (Zhang et al., 2020a; Raffel et al., 2020; Jia and Diaz-Rodriguez, 2025) (yielding k-LLMmeans). In both variants, summaries are recomputed only at scheduled iterations, so the algorithms share the same summary-as-centroid mechanism and differ solely in the summarizer. While methods from (1) and (2) are standard in summarization, our periodic use of summary-derived centroids within the k-means loop is, to our knowledge, novel and central to the method’s interpretability.
@@
Each summarization step produces a concise, human-readable centroid that exposes evolving cluster semantics, simplifying debugging, validation, and labeling. The summarizer is modular: it can be LLM-free (k-NLPmeans) or LLM-assisted under a fixed per-iteration budget that does not grow with dataset size (k-LLMmeans). For large-scale and streaming settings, we adapt mini-batch k-means (Sculley, 2010) by inserting summarization steps into the mini-batch update rule, enabling efficient online clustering while retaining interpretability.
@@
Recent work shows that LLMs can deliver strong unsupervised clustering (Zhang et al., 2023; Feng et al., 2024; De Raedt et al., 2023; Viswanathan et al., 2024; Shi and Sakai, 2023; Tarekegn et al., 2024; Nakshatri et al., 2023). However, many pipelines face two practical issues: (i) scalability, as they may rely on semi-supervision, iterative labeling, or a number of LLM calls that grows with dataset size; and (ii) opaque optimization, since they often combine prompts, greedy merges, and similarity thresholds without an explicit objective, making convergence behavior hard to analyze. Our k-LLMmeans addresses both: by summarizing per cluster at spaced iterations, we cap LLM usage independently of dataset size, and by keeping assignments and numeric updates in embedding space, we preserve the standard k-means objective between summary steps. In both k-NLPmeans and k-LLMmeans, if the summaries are weak, the procedure gracefully degrades to vanilla k-means with a suboptimal initialization, retaining its standard local-convergence behavior. Note that our method is LLM-optional: k-NLPmeans replaces the summarizer with classical NLP techniques, incurring no LLM usage.
@@ 2 Preliminaries: k-means for text clustering
$\min_{C_{1},C_{2},\dots,C_{k}}\sum_{j=1}^{k}\sum_{i\in[C_{j}]}\|\mathbf{x}_{i}-\bm{\mu}_{j}\|^{2},$ (1)
where $C_{j}$ denotes the set of embeddings assigned to cluster $j$, $C_{j}]=\{i|\mathbf{x}_{i}\in C_{j}\}$ denotes the set of embedding indices assigned to cluster $j$ and $\bm{\mu}_{j}$ is the cluster centroid, computed as the mean of the assigned embeddings:
$\bm{\mu}_{j}=\frac{1}{|C_{j}|}\sum_{i\in[C_{j}]}\mathbf{x}_{i}.$ (2)
Lloyd’s algorithm (Lloyd, 1982), the standard heuristic for k-means, alternates between assigning each document embedding $\mathbf{x}_{i}$ to its nearest centroid and recomputing each centroid as the mean of its assigned points. Repeating these two steps for $T$ iterations monotonically reduces the within‑cluster sum of squared distances, steering the procedure toward a (locally) optimal set of centroids. However, due to its sensitivity to initialization and the non-convex nature of its objective function, k-means does not guarantee convergence to the global optimum and can instead become trapped in local optima (MacQueen, 1967; Lloyd, 1982). Various strategies, such as k-means++ initialization and multiple restarts, have been proposed to mitigate these issues and improve the likelihood of achieving better clustering results (Arthur and Vassilvitskii, 2006).
@@
We enhance k-means for text clustering by periodically replacing the numerical centroid update with summarization steps that yield textual-summary–based centroids. The procedure (Algorithm 1 in Appendix D) is identical to k-means algorithm except that every $l$ iterations the mean update in equation 2 is replaced by the embedding of a cluster textual summary. At all other iterations, the standard update is used (see Figure 1 for an illustration with one summarization step). The summarizer can be instantiated with classical, deterministic methods (k-NLPmeans) or with LLM-based summaries (k-LLMmeans), and is not restricted to the specific techniques we discuss, any textual summarization operator can be dropped in.
@@
Our first variant uses classical extractive summarization to compute a textual prototype in place of the standard centroid update. Formally, we replace equation 2 with
$\bm{\mu}_{j}=\text{Embedding}\!\left(f^{(q)}_{\text{NLP}}(S_{j})\right),$ (3)
@@
  • Centroid-based summarization (Radev et al., 2004): compute the centroid of the sentence embeddings for $S_{j}$; rank sentences by cosine similarity to this centroid; concatenate the top $q$ sentences (by rank) to form $f^{(q)}_{\text{NLP}}(S_{j})$.
  • Graph-based (TextRank) (Mihalcea and Tarau, 2004): build a graph whose nodes are sentences in $S_{j}$ with edge weights given by pairwise cosine similarity of their embeddings; run a PageRank-style algorithm to score sentences; select and concatenate the top $q$.
  • LSA-style SVD in embedding space (Deerwester et al., 1990): stack the sentence embeddings for $S_{j}$, apply singular value decomposition, score sentences by contribution to the leading components, and concatenate the top $q$.
The resulting summary text $f^{(q)}_{\text{NLP}}(S_{j})$ is then embedded in the same vector space as the documents to yield the new centroid $\bm{\mu}_{j}$.
@@
If instead we generate the summary with an LLM, we have k-LLMmeans. Formally, we replace equation 2 by
$\bm{\mu}_{j}=\text{Embedding}(f_{\text{LLM}}(p_{j}))\\$ (4)
where $p_{j}=\text{Prompt}\left(I,\{d_{z_{i}}|z_{i}\sim[C_{j}]\}_{i=1}^{m_{j}}\right)$, and $m_{j}=\min(m,|C_{j}|)$. Here, $z_{i}\sim[C_{j}]$ denotes a sampled index of the embeddings assigned to cluster $C_{j}$ (without repetitions) and $m$ is a parameter that represents the maximum number of sampled indices used to compute the cluster centroid $\bm{\mu}_{j}$. In simple terms, we update a cluster’s centroid by using the embedding of the response generated by an LLM when queried with a prompt containing a summarization instruction $I$ and a representative sample of documents from the cluster. Rather than providing all documents within the cluster as input, the LLM processes a representative sample as a context prompt. While incorporating the entire cluster is theoretically possible, it poses practical challenges due to prompt length limitations. Therefore, we propose selecting the sample cluster documents using a k-means++ sampling of the cluster embeddings. Our experiments demonstrate that this sampling process facilitates a more effective synthesis of the cluster’s content, leading to improved summaries and, consequently, more refined centroid updates. The instruction $I$ varies depending on the clustering task, but standard summarization prompts are generally sufficient. Figure 1 illustrates how k-LLMmeans with a single summarization step enhances the standard k-means algorithm.
@@ 3.3 Advantages of our approaches
Over advanced LLM-based clustering methods. Our approach offers three key advantages over more complex LLM-based clustering methods: (1) Optimization landscape. Grounded in the standard k-means objective, we inherit algorithmic convergence without relying on the fragile heuristics common in newer LLM‑driven methods. A poor summary simply leads the procedure back toward a regular k-means local optimum, whereas competing approaches hinge on stable, format‑specific LLM outputs. (2) Scalability. Unlike most state-of-the-art methods, whose LLM usage complexity grows with sample size (Feng et al., 2024; De Raedt et al., 2023) or require fine-tuning (Zhang et al., 2023). k-NLPmeans requires no LLM calls and k-LLMmeans only performs $k$ LLM calls per summarization step, with even few summarization steps yielding substantial performance gains. (See Section 5.1). (3) Interpretability. Replacing numeric centroids with textual summaries turns each prototype into a concise, human-readable synopsis; practitioners can track how cluster semantics evolve over time without post-hoc labeling. This transparency extends naturally to our mini-batch variant, enabling real-time monitoring in streaming scenarios (see Figure 2 and Section 5.1).
@@ 4 Mini-batch k-NLPmeans and k-LLMmeans
Mini-batch k-means (Sculley, 2010) is an efficient strategy for large-scale text clustering that processes small, randomly sampled mini-batches instead of the full dataset. This approach substantially reduces memory usage and computational cost, making it well suited for continuously generated text streams, such as those from social media, news, or customer feedback; where data must be clustered incrementally without full dataset access. Mini-batch k-means exhibits convergence properties comparable to standard k-means while offering superior scalability.
Although numerous streaming clustering methods that do not rely on LLMs have been studied (Silva et al., 2013; Aggarwal, 2018; Ribeiro et al., 2017; Aggarwal et al., 2003; Ackermann et al., 2012; Ordonez, 2003), only a few incorporate LLMs (Tarekegn et al., 2024; Nakshatri et al., 2023). Moreover, existing offline LLM-based clustering approaches face scalability issues, highlighting the need for scalable summary-based clustering in an online setting. To address this, we introduce mini-batch k-NLPmeans and k-LLMmeans, which directly extend mini-batch k-means by inserting summarization steps into the mini-batch update.
@@ 5 Static experiments
We use four benchmark datasets that span diverse domains and classification granularities: Bank77 (Casanueva et al., 2020), CLINC (Larson et al., 2019), GoEmo (Demszky et al., 2020) and MASSIVE (domain and intent) (FitzGerald et al., 2023). See Appendix A.1 for detailed descriptions of the datasets. We evaluate our algorithms on each of the four datasets using the known number of clusters and performing 120 centroid-update iterations. We calculate two variants of our algorithms differing in the number of summarization steps. The single variant uses a single summarization step ($l=60$), while the multiple variant performs five summarization steps ($l=20$). To demonstrate the robustness of our approach, we compute embeddings with models: DistilBERT(Sanh, 2019), e5-large(Wang et al., 2022), S-BERT(Reimers and Gurevych, 2019), and text-embedding-3-small(OpenAI, 2024). For k-NLPmeans we evaluate using TextRank, Centroid and LSA summarization methods mentioned in Section 3.1, with $q=5$. For the LLM component of k-LLMmeans, we use GPT-3.5-turbo(OpenAI, 2023), GPT-4o(Hurst et al., 2024), Llama-3.3(Grattafiori et al., 2024), Claude-3.7(Anthropic, 2025) and DeepSeek-V3(Liu et al., 2024). For the instruction task $I$, we employ a simple summarization prompt depending on the task. For example, for Bank77 we use the prompt: “The following is a cluster of online banking questions. Write a single question that represents the cluster concisely.” We explore the effect of prompt size by summarizing using all cluster documents as input, and a few-shot (FS) variant where each prompt includes only $m=10$ randomly selected documents instead of the full cluster. For each setting we run five different seeds.
As traditional clustering baselines, we consider k-means, k-medoids (Kaufman and Rousseeuw, 2008), spectral clustering (Ng et al., 2001), agglomerative clustering (Johnson, 1967), and Gaussian Mixture Models (GMM) (Dempster et al., 1977). We use ground-truth number of clusters and ten different seeds. We also include BERTopic (Grootendorst, 2022) as a strong topic-modeling baseline. In addition, we report results obtained by applying our proposed methods to the document embeddings produced by BERTopic. For advanced LLM-based methods, we compare with results obtained by Feng et al. (2024) on ClusterLLM (Zhang et al., 2023), IDAS (De Raedt et al., 2023), and LLMEdgeRefine (Feng et al., 2024). All k-means based methods are initialized with k-means++.